# Differential Privacy of Cross-Attention with Provable Guarantee

### v1

@@ 4.2 Softmax Activation
In this section, we present the algorithm (Algorithm 3) that can answer the weighted softmax query (Definition 4.1) and can be further used to design DP cross-attention. First, we introduce the definition of weighted softmax query, which is also an abstraction for the problem described in Section 3.4.
Definition 4.1 (Softmax query).
Let $\phi$ be the distance measurement function defined in Definition 1.1. For the dataset $X\in[0,R]^{n\times d}$ where $x_{i}^{\top}$ is the $i$-th row of $X$ and query $y\in[0,R]^{d}$, we define the weighted exponential inner product/softmax query to be:
$\sum_{i\in[n]}w_{i}\phi(x_{i},y)=\sum_{i\in[n]}w_{i}\exp(\langle x_{i},y\rangle/d)=w^{\top}\exp(Xy/d).$
Based on the Definition 4.1, we design a novel algorithm that can make weighted softmax query DP. More details can be found in Section G.
Theorem 4.2 (Softmax query, informal version of Theorem G.8).
Let $R\geq 1$. Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$. Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition G.3). Let the accuracy parameter be $\epsilon_{s}\in(0,0.1)$. There is a data structure DPTreeSoftmax (Algorithm 3) that uses $O(nr)$ spaces to solve softmax query problem for dataset $X\subset[0,R]^{d}$ and support the following operations:
@@
    • $z$ is $(\epsilon,\delta+\delta^{\prime})$-DP private of $w^{\top}\exp(Xy/d)$,
    • $|z-w^{\top}\exp(Xy/d)|\leq(\alpha+\epsilon_{s})\cdot w^{\top}\exp(Xy/d)+~{}O(\epsilon^{-1}\alpha^{-1/2}\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(1/\delta^{\prime})}\cdot\log^{3/2}n)$,
    • it holds with probability $0.99$.
@@
In Theorem 4.2, the parameter $\epsilon_{s}$ is the accuracy parameter for polynomial kernel approximation described in Section 6.5. Besides, note that the error bound in Theorem 4.2 does not depend on $\delta$ but depends on $\delta^{\prime}$. The role of $\delta$ is to control a hidden constant term in the big $O$ notation, i.e., increasing $\delta$ can reduce the error by a small constant (Fact 3.7). In practice, we set $\delta$ as a small positive constant close to 00. Please refer to the Lemma B.6 for more details.
@@ 4.3 Adaptive Query Data Structure
We can now present the theorem for our data structure that can answer weighted softmax queries with a DP guarantee and concrete error bound, and it is robust to adaptive queries. The corresponding proofs can be found in Section F and Section G.
Theorem 4.4 (Adaptive query softmax data structure, informal version of Theorem G.11).
Let $R\geq 1$. Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$. Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition G.3). Let the accuracy parameter be $\epsilon_{s}\in(0,0.1)$. Let $X\in[0,R]^{n\times d}$ be the dataset, $w\in[-R_{w},R_{w}]^{n}$ be weights, $y\in[0,R]^{d}$ be the query, $\alpha\in(0,1)$ be the relative error parameter and $p_{f}$ be the failure probability parameter. Let $l=O(r\log(dR/(\epsilon_{s}p_{f})))$. There is a data structure DPTreeSoftmaxAdaptive (Algorithm 4) that uses $O(lnr)$ spaces to solve the weighted softmax query problem for the dataset $X\subset[0,R]^{d}$ and supports the following operations:
  • Init$(X\subset[0,R]^{d},n\in\mathbb{N}_{+},w\in[-R_{w},R_{w}]^{n},\epsilon\in(0,1),\delta\in(0,1),\delta^{\prime}\in(0,1),c\in(0,0.1),\epsilon_{s}\in(0,0.1),p_{f}\in(0,0.01))$. (Algorithm 4) It takes $O(lnr)$ time to initialize the data structure.
  • DistanceQuery$(y\in[0,R]^{d},\alpha\in(0,1))$. (Algorithm 4) It takes $O(\alpha^{-1}lr\log^{2}n)$ time to output a number $z$ such that
    • $z$ is $(\epsilon,\delta+\delta^{\prime})$-DP private of $w^{\top}\exp(Xy/d)$,
    • $|z-w^{\top}\exp(Xy/d)|\leq(\alpha+\epsilon_{s})\cdot w^{\top}\exp(Xy/d)+~{}O(\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$,
    • it holds with probability $1-p_{f}$ (where $p_{f}$ is used in $l$),
@@
We describe the parallelization of our algorithms. In the second for loop of Init and the for loop of DistanceQuery in Algorithm 3, the $r$ DPTreeDistance data structures instantiated for each coordinate are independent of each other. In addition, the for loops in Algorithm 4 can also be parallelized since the $l=O(r\log(dR/(\epsilon_{s}p_{f})))$ copies are independent. After parallelization, we have the final time complexity of Init to be $O(nr)$ and DistanceQuery to be $O(\alpha^{-1}\log^{2}n)$ in Algorithm 4 with $O(lr)$ GPU process.
In Theorem 4.4, we show that our DPTreeSoftmaxAdaptive data structure defined above have $(\epsilon,\delta)$-DP guarantee with $\alpha$ relative error and $O(\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}(n)+w^{\top}\exp(Xy/d))$ additive error under $O(lnr)$ memory consumption, $O(lnr)$ initialization time and $O(\alpha^{-1}lr\log^{2}n)$ running time under adaptive query. We can apply our data structure in cross-attention, and we show detail in the following section.
5 Main Results: Cross-Attention
In this section, we show our main result for cross-attention. Theorem 5.1 states that we can make the cross-attention module DP.
Theorem 5.1 (Cross-attention, informal version of Theorem G.12).
Let $Q,K,V,\phi,\mathrm{Attn}$ be defined in Definition 1.1. Let $\alpha\in(0,1)$ be the relative error parameter and $p_{f}$ be the probability of failure parameter. Let $r,s,\epsilon_{s}$ be the parameters of the polynomial kernel methods (Lemma 6.7). Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition G.3). Let $l=O(r\log(dR/(\epsilon_{s}p_{f})))$. Let $\phi$ be softmax function (Definition 4.1). There is a data structure DPTreeSoftmaxAdaptive (Algorithm 4) that uses $O(lnrd)$ spaces to make cross-attention DP and supports the following operations:
    • $z$ is $(\epsilon,\delta+\delta^{\prime})$-DP,
    • $z$ has relative error $n^{-1}(\alpha+\epsilon_{s})$,
    • $z$ has additive error $O(n^{-1}\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$,
    • it holds with probability $1-p_{f}$ (where $p_{f}$ is used in $l$),
    • it is robust to adaptive query.
In Theorem 5.1, we use our DPTreeSoftmaxAdaptive (see Section 6.6) and guarantee that for each query token of cross-attention, the output is $(\epsilon,\delta)$-DP with $n^{-1}(\alpha+\epsilon_{s})$ relative error and $O(n^{-1}\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$ additive error, and $O(\alpha^{-1}lr\log^{2}n)$ running time under adaptive query. More specifically, the algorithm creates $d$ DPTreeSoftmaxAdaptive data structures, each requiring $O(lnr)$ memory consumption and $O(lnr)$ initialization time.
Thus, our algorithm can theoretically protect system prompts/RAG data in cross-attention as discussed in Section 1. In Section 6, we provide a detailed technical overview, and in Section 7, we will present self-attention and DP-related discussion.
6 Technique Overview

### v2

@@ 2.2 Differential Privacy Definitions
3 Main Results: Cross-Attention
In this section, we show our main result for cross-attention. Theorem 3.1 states that we can ensure the entire cross-attention module satisfies DP and is robust to adaptive queries. Our high-level idea is based on the similarity between weighted distance problem and cross-attention. For a typical weighted distance problem, we define the following: Let $w\in\mathbb{R}^{n}$ be the weights, $X\in\mathbb{R}^{n\times d}$ be the data matrix, where $x_{i}^{\top}$ is the $i$-th row of $X$ for $i\in[n]$, and let $y\in\mathbb{R}^{d}$ be the query. Suppose we need to answer $\ell_{1}$-distance query. We have
$\sum_{i\in[n]}\underbrace{w_{i}}_{\mathrm{weight}}\|\underbrace{y}_{\mathrm{query}}-\underbrace{x_{i}}_{\mathrm{data}}\|_{1}.$
[…]
$(AV)_{i,k}=$ $~{}\sum_{j=1}^{n}\underbrace{V_{j,k}}_{\mathrm{weight}}\exp(\langle\underbrace{Q_{i}}_{\mathrm{query}},\underbrace{K_{j}}_{\mathrm{data}}\rangle/d),$ (1)
which can be viewed as a weighted Softmax problem, where $V$ provides the weights, $Q$ is the query, and $K$ is the dataset. Thus, we choose to add noise to $K$ and $V$ based on the similarity between the weighted distance problem and cross-attention. Furthermore, we find that we can only handle one column of $V$, i.e., $V_{*,k}\in\mathbb{R}^{n}$, in a single data structure. Therefore, we need to initialize a total of $d$ different data structures, each with weights $V_{*,k}$ for $k\in[d]$.
Here, we present our main result below.
Theorem 3.1 (Softmax cross-attention, informal version of Theorem I.12).
Let $Q,K,V,\mathrm{Attn}$ be defined in Definition 1.1. Let $\alpha\in(0,1)$ be the relative error parameter and $p_{f}$ be the probability of failure parameter. Let $r,s,\epsilon_{s}$ be the parameters of the polynomial kernel methods (Lemma C.7). Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition I.3). Let $l=O(r\log(dR/(\epsilon_{s}p_{f})))$. There is a data structure DPTreeSoftmaxAdaptive (Algorithm 1) that uses $O(lnrd)$ spaces to ensure cross-attention satisfies DP and supports the following operations:
    • the process of output $z$ satisfies $(\epsilon,\delta+\delta^{\prime})$-DP,
    • the process of output $z$ has relative error $n^{-1}(\alpha+\epsilon_{s})$,
    • the process of output $z$ has additive error $O(n^{-1}\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$,
    • it holds with probability $1-p_{f}$ (where $p_{f}$ is used in $l$),
    • it is robust to adaptive query.
In Theorem 3.1, we use our DPTreeSoftmaxAdaptive (Algorithm 1) and guarantee that, for each query token of cross-attention, the output process satisfies $(\epsilon,\delta)$-DP with $n^{-1}(\alpha+\epsilon_{s})$ relative error and $O(n^{-1}\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$ additive error, and $O(\alpha^{-1}lr\log^{2}n)$ running time under adaptive query. More specifically, the algorithm creates $d$ DPTreeSoftmaxAdaptive data structures, each requiring $O(lnr)$ memory consumption and $O(lnr)$ initialization time. Notably, our error is inversely proportional to $n$, meaning that as the input token length increases, both the relative and approximate errors approach zero. This is achieved by the normalizing matrix $D$ (Definition 1.1). We refer the reader to Section I for proof details.
Thus, our algorithm theoretically protects system prompts/RAG data in cross-attention as discussed in Section 1. In Section 4, we provide a detailed technical overview, and in Section 5, we will present self-attention and DP-related discussion.
Algorithm 2 DPTree initialization and query
1:datastructure DPTree $\triangleright$ Theorem C.1
2:members
3:     $b:\mathbb{R}^{2n-1}$
4:     $c:\mathbb{R}^{2n-1}$
5:end members
6:procedure Init($a\in\mathbb{R}^{n},n\in\mathbb{N}_{+},\Delta\in\mathbb{R},\epsilon\in(0,1),\delta\in(0,1)$) $\triangleright$ Lemma D.3, Lemma C.3
7:     $b[n,2n-1]\leftarrow a$
8:     for $i=n\to 2n-1$ do
9:         $c[i]\leftarrow b[i]+\mathrm{TLap}(\Delta,\epsilon/\log n,\delta/\log n)$
10:     end for
11:     for $i=(\log n)\to 1$ do
12:         for $j=1\to 2^{i-1}$ do
13:              $k\leftarrow 2^{i-1}+j-1$
14:              $b[k]\leftarrow b[2k]+b[2k+1]$
15:              $c[k]\leftarrow b[k]+\mathrm{TLap}(\Delta,\epsilon/\log n,\delta/\log n)$
16:         end for
17:     end for
18:end procedure
19:procedure Query($x\in[n],y\in[n]$) $\triangleright$ Lemma D.4, D.5, D.6
20:     Trace from bottom nodes of $x$ and $y$ to find their lowest common ancestor, then we report the summation (based on $c$) by using at most $2\log n$ nodes on the path. Let Value be the above summation.
21:     return Value
22:end procedure
23:procedure TrueQuery($x\in[n],y\in[n]$)
24:     Trace from bottom nodes of $x$ and $y$ to find their lowest common ancestor, then we report the summation (based on $b$) by using at most $2\log n$ nodes on the path, where the height of the tree is $\log n$, and we need left and right boundary points. Let Value be the above summation.
25:     return Value
26:end procedure
27:end datastructure
4 Key Data Structure: DPTree
This section provides our key data structures: DPTree (Algorithm 2), DPTreeDistance (Algorithm 5 and 6), DPTreeHighDim (Algorithm 7), DPTreeSoftmax (Algorithm 3), and DPTreeSoftmaxAdaptive (Algorithm 1).
In Section 4.1, we provide our high-level proof insights. In Section 4.2, we give our basic building block algorithms DPTree, DPTreeDistance and DPTreeHighDim. In Section 4.3, we present our DPTreeSoftmax algorithm that solves the weighted Softmax problem. In Section 4.4, we present our DPTreeSoftmaxAdaptive algorithm that enables DPTreeSoftmax to handle adaptive query problem.
4.1 Technique Overview
Notice that Eq. (1) is not a typical distance measure like $\ell_{1}$ or $\ell_{2}$, but by using polynomial kernel method techniques, we transform it into a distance measure. AS [6] states that the exponential inner product can be approximated by polynomial kernel function $P(\cdot):\mathbb{R}^{d}\to\mathbb{R}^{r}$, i.e., $P(x)^{\top}P(y)\approx\exp(x^{\top}y/d)$ for two vector $x,y\in\mathbb{R}^{d}$, with a relative error. Then, by the Law of Cosines, we transform the inner product of polynomial kernel functions into a distance measure, i.e.,
$2P(x)^{\top}P(y)=$ $~{}-\|P(x)-P(y)\|_{2}^{2}+\|P(x)\|^{2}_{2}+\|P(y)\|^{2}_{2}.$ (2)
After transforming Eq. (1) into a distance measure, we design the DPTree series data structures to provide cross-attention DP guarantee.
In summary, we first design the data structure DPTree (Algorithm 2) that builds a binary segment tree with truncated Laplace noise added in the leaf nodes to ensure DP guarantee. Then, based on this data structure, we design DPTreeDistance (Algorithm 5 and 6) to answer one dimensional weighted distance queries $\sum_{i=1}^{n}w_{i}\cdot|y-x_{i}|$, which utilizes DPTree to store and return noised weights $w_{i}$ multiplied with the approximated distances between the query $y$ and data $x_{i}$. We further decompose high dimensional $\ell_{p}^{p}$-distance problem into one dimensional $\ell_{1}$-distance problems using
$\sum_{i=1}^{n}w_{i}\cdot\|y-x_{i}\|_{p}^{p}=$ $~{}\sum_{k=1}^{d}\sum_{i=1}^{n}w_{i}\cdot|y_{k}-x_{i,k}|^{p}.$ (3)
Based on this decomposition, we design DPTreeHighDim (Algorithm 7) which is capable of answering high dimension queries. Then, using Eq. (2) and DPTreeHighDim, we design DPTreeSoftmax (Algorithm 3) to answer Softmax queries. By building multiple copies of this data structure, we boost the success probability such that it can answer any query (including adaptive query) with an additive error, establishing the final data structure DPTreeSoftmaxAdaptive (Algorithm 1). See Section C for a more detailed outline of algorithms and proof techniques.
4.2 DPTree, DPTreeDistance, and DPTreeHighDim
We design a basic data structure DPTree (Algorithm 2) that answers summation queries by a summation segment tree with truncated Laplace noise (Definition 2.4). The algorithm first builds a binary summation tree in an array and then adds truncated Laplace noises to each node. In query time, we first trace from bottom nodes to find their lowest common ancestor, then report the summation by using at most $2\log n$ nodes on the path (Algorithm 2). Based on the parallel composition rule of DP (Fact B.7), we find that if we have multiple disjoint interval queries, the error of the weighted sum of the intervals can be bounded independently of the number of queries (Lemma D.8). See more details in Section D.
We then design DPTreeDistance, a one-dimensional weighted $\ell_{1}$ distance data structure detailed in Algorithm 5 and 6. Initialization involves rounding each data point to the nearest multiple of a small interval and aggregating their weights into an array (illustrated in Figure 1), which is then input into our DPTree. At query time, we retrieve aggregated weights within small intervals and multiply them by their distances to the query point. We introduce a relative error parameter $\alpha$ to reduce the number of iterations to $O(\log(n)/\alpha)$, improving efficiency. Guided by Eq.(3), we design DPTreeHighDim (Algorithm 7), which extends DPTreeDistance to higher dimension by constructing independent data structures for each coordinate. See details in Section F and G.
Algorithm 3 Softmax query
1:datastrucutre DPTreeSoftmax $\triangleright$ Theorem 4.2
2:members
3:     ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{r}:\textsc{DPTreeDistance}$triangleright Algorithm 5, Theorem I.7
4:     $P:[0,\Gamma_{R,s}]^{n\times r}$triangleright Definition I.3 for $\Gamma_{R,s}$, Eq. (9) for $s$, Eq. (10) for $r$
5:     $w:[-R_{w},R_{w}]^{n}$
6:     $P_{wx},~{}s_{w},~{}\epsilon_{s}:\mathbb{R}$
7:end members
8:procedure Init($X\subset[0,R]^{d}$, $n\in\mathbb{N}_{+}$, $w\in[-R_{w},R_{w}]^{n}$, $\epsilon\in(0,1)$, $\delta\in(0,1)$, $\delta^{\prime}\in(0,1)$, $c\in(0,0.1)$, $\epsilon_{s}\in(0,0.1)$) $\triangleright$ Lemma C.7
10:     for $j=1\to n$ do
11:         Compute $P(x_{j})$triangleright Polynomial kernel function $P(\cdot)$, Lemma I.5
12:         Compute $w_{j}\|P(x_{j})\|_{2}^{2}$
13:         $P_{wx}\leftarrow P_{wx}+w_{j}\|P(x_{j})\|_{2}^{2}$
14:         $s_{w}\leftarrow s_{w}+w_{j}$
15:         $P_{j,:}\leftarrow P(x_{j})$
16:     end for
17:     for $i=1\to r$ do
18:         $\mathcal{D}_{i}.$Init($P_{:,i}$, $n$, $w$, $c\epsilon/\sqrt{r\log(1/\delta^{\prime})}$, $\delta/r$) $\triangleright$ Algorithm 5
19:     end for
20:end procedure
21:procedure DistanceQuery($y\in[0,R]^{d}$, $\alpha\in(0,1)$) $\triangleright$ Lemma C.7
22:     Value $\leftarrow 0$
23:     Compute $P(y)$
24:     Compute $\|P(y)\|_{2}^{2}$
25:     for $i=1\to r$ do
26:         Value $\leftarrow$ Value + $\mathcal{D}_{i}$.DistanceQuery$(P(y)_{i},\alpha)$triangleright Algorithm 6
27:     end for
29:     return Value
30:end procedure
31:end datastrucutre
@@ 4.3 Softmax Activation
In this section, we present DPTreeSoftmax (Algorithm 3) that answers the weighted Softmax query (Definition 4.1) and is further used to design DP cross-attention. First, we introduce the definition of weighted Softmax query, an abstraction for the problem described in Eq. (1).
Definition 4.1 (Weighted Softmax query (without normalization)).
For the dataset $X\in[0,R]^{n\times d}$ where $x_{i}^{\top}$ is the $i$-th row of $X$ and query $y\in[0,R]^{d}$, we define the weighted exponential inner product/Softmax query to be:
$\sum_{i\in[n]}w_{i}\exp(\langle x_{i},y\rangle/d)=w^{\top}\exp(Xy/d).$
Building on Definition 4.1, we develop a novel algorithm to answer differentially private weighted Softmax queries using the polynomial kernel method from AS [6]. Specifically, in Eq.(2), there is a term that computes the weighted $\ell_{2}^{2}$ distance, which we calculate using DPTreeHighDim. We then compute the exact term for the weighted $\ell_{2}^{2}$ norms of the approximation kernel. By summing these terms with a controlled error, we extend DPTreeHighDim to answer the Softmax query efficiently. More details can be found in Section I.
Theorem 4.2 (Softmax query, informal version of Theorem I.8).
Let $R\geq 1$. Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$. Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition I.3). Let the accuracy parameter be $\epsilon_{s}\in(0,0.1)$. Our data structure DPTreeSoftmax (Algorithm 3) uses $O(nr)$ spaces to solve Softmax query problem for dataset $X\subset[0,R]^{d}$ and support following operations:
@@
    • the process of output $z$ satisfies $(\epsilon,\delta+\delta^{\prime})$-DP private, which computes $w^{\top}\exp(Xy/d)$,
    • the error bound satisfies $|z-w^{\top}\exp(Xy/d)|\leq(\alpha+\epsilon_{s})\cdot w^{\top}\exp(Xy/d)$
      $+~{}O(\epsilon^{-1}\alpha^{-1/2}\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(1/\delta^{\prime})}\cdot\log^{3/2}n)$,
    • it holds with probability at least $0.99$.
@@
In Theorem 4.2, the parameter $\epsilon_{s}$ is the accuracy parameter for polynomial kernel approximation described in Section C.5. Besides, note that the error bound in Theorem 4.2 does not depend on $\delta$ but depends on $\delta^{\prime}$. The role of $\delta$ is to control a hidden constant term in the big $O$ notation, i.e., increasing $\delta$ reduces the error by a small constant (Fact 2.5). In practice, we set $\delta$ as a small positive constant close to 00. Please refer to the Lemma D.6 for more details.
@@ 4.4 Adaptive Query Data Structure
We adapt our DPTreeSoftmax to DPTreeSoftmaxAdaptive (Algorithm 1) to solve the adaptive query problem. By proving it can handle any query within the query space with a certain error, we ensure it effectively processes adaptive queries. We first boost the constant probability to high probability using the Chernoff bound (Lemma B.2). Employing an $\epsilon_{0}$-net argument and the union bound, we bound all query points within the net. Finally, we use the Lipschitz property of the weighted Softmax distance function with an additive error to bound all points in the query space. The corresponding proofs can be found in Section H and Section I.
Theorem 4.4 (Adaptive query Softmax data structure, informal version of Theorem I.11).
Let $R\geq 1$. Let $r\leq{2s+2d\choose 2s}$ and $s=O(\max\{\frac{\log(1/\epsilon_{s})}{\log(\log(1/\epsilon_{s})/R)},R^{2}\})$. Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition I.3). Let the accuracy parameter be $\epsilon_{s}\in(0,0.1)$. Let $X\in[0,R]^{n\times d}$ be the dataset, $w\in[-R_{w},R_{w}]^{n}$ be weights, $y\in[0,R]^{d}$ be the query, $\alpha\in(0,1)$ be the relative error parameter and $p_{f}$ be the failure probability parameter. Let $l=O(r\log(dR/(\epsilon_{s}p_{f})))$. There is a data structure DPTreeSoftmaxAdaptive (Algorithm 1) that uses $O(lnr)$ spaces to solve the weighted Softmax query problem for the dataset $X\subset[0,R]^{d}$ and supports the following operations:
  • Init$(X\subset[0,R]^{d},n\in\mathbb{N}_{+},w\in[-R_{w},R_{w}]^{n},\epsilon\in(0,1),\delta\in(0,1),\delta^{\prime}\in(0,1),c\in(0,0.1),\epsilon_{s}\in(0,0.1),p_{f}\in(0,0.01))$. It takes $O(lnr)$ time to initialize the data structure.
  • DistanceQuery$(y\in[0,R]^{d},\alpha\in(0,1))$. It takes $O(\alpha^{-1}lr\log^{2}n)$ time to output a number $z$ such that
    • the process of output $z$ satisfies $(\epsilon,\delta+\delta^{\prime})$-DP private, which computes $w^{\top}\exp(Xy/d)$,
    • the error bound satisfies $|z-w^{\top}\exp(Xy/d)|\leq(\alpha+\epsilon_{s})\cdot w^{\top}\exp(Xy/d)$
      $+~{}O(\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$,
    • it holds with probability at least $1-p_{f}$ (where $p_{f}$ is used in $l$),
@@
We describe the parallelization of our algorithms. In the second for loop of Init and the for loop of DistanceQuery in Algorithm 3, the $r$ DPTreeDistance data structures instantiated for each coordinate are independent of each other. In addition, the for loops in Algorithm 1 are also parallelizable since the $l=O(r\log(dR/(\epsilon_{s}p_{f})))$ copies are independent. After parallelization, we have the final time complexity of Init to be $O(nr)$ and DistanceQuery to be $O(\alpha^{-1}\log^{2}n)$ in Algorithm 1 with $O(lr)$ GPU process.
5 Discussion
How do we extend to self-attention?
Why not add noise to some other places?
Where and how to add DP noises is an important problem to ask during the DP algorithm design. In this paper, we consider the problem of $\sum_{i=1}^{n}w_{i}\exp(\langle x_{i},y\rangle/d)$ where $y,x_{i}\in[0,R]^{d}$ and $w\in[-R_{w},R_{w}]^{n}$ (Definition 4.1). Notice that the only place where we add noises is in the most basic building block data structure DPTree (Algorihtm 2). From Lemma C.3 and the way we initialize DPTree in Algorithm 5, we see that the sensitivity $\Delta$ of this problem is $2R_{w}$.
A simple method for adding noise involves adding $n$ noises to a length $n$ array, with each item $w_{i}\exp(\langle x_{i},y\rangle/d)$ for $i\in[n]$. However, this approach increases the error by a factor of $n$ by basic composition (Fact B.6) and also makes the model dependent on the number of queries. Besides, it only supports a single query and requires rebuilding the tree for each new query, rendering it impractical. In contrast, our current noise-adding technique (Lines 9 and 15 of Algorithm 2) utilizes a summation tree such that the error only increases by a factor of $\operatorname{poly}\log n$. This method also supports multiple queries, eliminating the need to rebuild the tree each time.
How to remove the relative error parameter $\alpha$?
The relative error parameter $\alpha$ in Theorem 3.1 appears because of the $(1+\alpha)$-approximation introduced in Algorithm 5 (Remark F.3) to reduce the number of required iterations from naive $O(n)$ to $O(\log(n)/\alpha)$. However, we notice that a recent work [59] does not utilize $(1+\alpha)$-approximation and still achieves $O(\log n)$ iteration number. They introduce a new tree node representation where each node stores the sum of distances from one point to multiple points, enabling the answer to be divided into only $\log n$ values, each combining two distance values, two count values, and $y$ itself. Our DPTree algorithms can be integrated with their method, thus removing parameter $\alpha$.
6 Conclusion
Acknowledgement
Appendix
Roadmap.
The appendix is organized as follows. In Section A, we provide more related works. In Section B, we give the preliminary of our paper. In Section C, we offer an outline of our proof techniques. In Section D, we give the analysis of the data structure DPtree that can solve summation problem with DP and accuracy guarantee. In Section E, we show how to solve weighted distance problem. In Section F, we give our DPTreeDistance data structure that can solve one dimensional $\ell_{1}$ distance problem with DP and accuracy guarantee. In Section G, we present the analysis of our DPTreeHighDim (Algorithm 7) data structure, which can address the high-dimensional $\ell_{1}$ distance problem while ensuring differential privacy and accuracy guarantees. In Section H, we show how we can handle adaptive query. In Section I, we show how to extend our algorithm to Softmax activation and give the analysis of DPTreeSoftmax (Algorithm 3) and DPTreeSoftmaxAdaptive (Algorithm 1).
Appendix A More Related Work
Differential Privacy Guarantee Analysis.
Ever since DMNS [19] proposes the notion of differential privacy (DP), it has become one of the most essential standards of privacy protection in both theoretical and empirical ways [22, 67, 115, 84, 111]. DP provides a powerful, robust, and quantifiable privacy definition, allowing algorithm design with concrete privacy and accuracy guarantee [50, 24, 3, 61, 51, 37, 10, 12, 25, 16, 46, 78, 79, 56, 62, 28, 29, 60, 17, 13, 18, 27, 40, 39, 64, 38, 23, 95, 21, 107, 94, 42, 70, 66]. Additionally, new mechanisms have been proposed beyond the traditional Laplace, Gaussian, and Exponential mechanisms [20]. For example, truncated Laplace mechanism [32] is proved to be the current tightest the lower and upper bounds on the minimum noise amplitude and power cross all $(\epsilon,\delta)$-DP distributions.
Cross-Attention in System Prompt, RAG, Stable Diffusion and More.
Cross-attention [97], first introduced in language translation, is a widely used technique in many advanced AI systems. For example, Stable Diffusion [87, 73, 98, 103, 106] and SORA [81] employ cross-attention as a core module for a text-to-image conditional generation. This technique is also utilized by other multimodal models [72], including Imagen [90] and Diffusion Transformer [85]. In the realm of text-to-image editing, HMT+ [49] analyzes and controls the cross-attention module to enable editing without requiring additional training. Furthermore, YYB+ [114] tackles the issue of inaccurate cross-attention maps, enhancing fine-grained control over edited regions while preventing unintended changes to other areas. In addition, Retrieval Augmented Generation (RAG) [68, 11, 43], a technique that improves model responses by retrieving information from a knowledge base or external documents, extensively uses cross-attention as its core design module. Cross-attention also has other applications. ORST [82] demonstrates that the prompt-tuning [71] task can be formulated as cross-attention, while CFP [14] uses cross-attention to fuse multi-scale features in vision transformers, thereby reducing computation. Moreover, attention-based Transformer architecture makes LGMs equipping many emergent ability [104], such as spatial reasoning [101], mathematical reasoning [65], in-context learning ability [93], compositional ability [109], few-shot adaptation ability [89, 110], and so on. There are some other works that used cross attention in Hopfield Models [52, 100, 48, 108, 99, 44, 45].
Appendix B More Preliminary
In Section B.1, we give the probability tools we use in the paper. In Section B.2, we provide the algebraic facts we use. In Section B.3, we give the DP facts we use in the paper. In Section B.4, we compare between popular DP mechanisms.
B.1 Probability Tools
In this section, we give several probability lemmas.
Lemma B.1 (Markov’s inequality).
If $x$ is a nonnegative random variable and $t>0$, we have
$\Pr[x\geq t]\leq\frac{\operatorname*{{\mathbb{E}}}[x]}{t}.$
Lemma B.2 (Chernoff bound, [15]).
$\Pr[x\geq(1+\delta)\mu]\leq\exp(-\delta^{2}\mu/3),$
and for all $10<\delta<$ 1
$\Pr[x\leq(1-\delta)\mu]\leq\exp(-\delta^{2}\mu/2).$
Lemma B.3 (Chebyshev’s inequality).
$\Pr[|x-\mu|\geq k\sigma]\leq\frac{1}{k^{2}}.$
B.2 Algebraic Facts
Fact B.4 (Upper bound of exponential, Fact C.9 in LSSZ24b [73]).
For $a\in\mathbb{R}$, $b\in\mathbb{R}$, $a,b\leq R$, where $R\geq 0$, we have
$|\exp(a)-\exp(b)|\leq\exp(R)|a-b|.$
B.3 DP Facts
Fact B.5 (Post-processing, see Fact 2.1 in GKK+ [37]).
Fact B.6 (Basic composition, see Fact 2.3 in GKK+ [37]).
We can do much better if we know that the inputs are disjoint.
Fact B.7 (Parallel composition, see Fact 2.4 in GKK+ [37]).
Theorem B.8 (Advanced composition, see Theorem 3.20 in DR [20]).
B.4 Comparison of Truncated Laplace, Gaussian, and Laplace Mechanisms
We first define the Laplace mechanism as below:
Definition B.9 (Laplace distribution).
We use $\mathrm{Lap}(b)$ to denote the pdf: $p(z)=\frac{1}{2b}\exp(-\frac{|z|}{b}).$
Fact B.10.
In this paper, we use the Chebyshev inequality to bound the error, and from GDGK [32], we know that the truncated Laplace mechanism has the current minimum variance across all $(\epsilon,\delta)$-DP distributions.
The variance of Gaussian mechanism in Theorem 3.22 in DR [20]:
The variance of Laplace mechanism in Fact B.10:
$\operatorname*{\mathrm{Var}}=\frac{2\Delta^{2}}{\epsilon^{2}}.$
The variance of truncated Laplace mechanism in Fact 2.5, for $c\in(0,1]$:
$\operatorname*{\mathrm{Var}}=\frac{2\Delta^{2}c}{\epsilon^{2}}.$
Appendix C Proof Outline